# Code Equivalence, Point Set Equivalence, and Polynomial Isomorphism

### v1

@@ 6. Doublings of Sets of Points
For a proof of (a), see [23, Kor. 6.13] or [7, Prop. 3.3.18]. Claim (b) follows from the Hilbert function of ${\mathfrak{J}}_{R/{\mathbb{F}}_{q}[\ell]}$ given in Section 5, and (c) follows immediately from (b). ∎
For further properties of doublings, see for instance [26, Sect. 2.5]. In the case of self-associated sets of points, we obtain the following corollary.
@@
The goal of this section is to reformulate the Point Set Equivalence Problem in several ways. A linear change of coordinates $\Lambda:\;{\mathbb{P}}^{k-1}\longrightarrow{\mathbb{P}}^{k-1}$ induces homomorphisms of several algebraic objects considered above in the follows ways.
@@
$\varPhi^{\prime}(\lambda_{\omega}({\varphi}))\;=\;c_{1}(\ell^{\prime})^{d}f^{\prime}_{1}+\cdots+c_{n}(\ell^{\prime})^{d}f^{\prime}_{n}\;=\;\lambda_{R}(\ell^{d}(c_{1}f_{1}+\cdots+c_{n}f_{n}))\;=\;\lambda_{R}(g)$
@@
Finally, we prove that (2) implies (1). We use $i=2$ and note that the initial degree of $\ell^{2}\,{\mathfrak{J}}_{R/{\mathbb{F}}_{q}[\ell]}$ is $r_{\mathbb{X}}+2$. Therefore the hypothesis that $\lambda_{D,2}$ is an isomorphism implies that $(\lambda_{R})_{d}:\;R_{d}\longrightarrow R^{\prime}_{d}$ is an isomorphism of ${\mathbb{F}}_{q}$-vector spaces for every $d\leq r_{\mathbb{X}}+1$. Consequently, we have $\lambda((I_{\mathbb{X}})_{d})=(I_{{\mathbb{X}}^{\prime}})_{d}$ for every $d\leq r_{\mathbb{X}}+1$. In particular, we see that $\mathbb{X}$} and $\mathbb{X}$$^{\prime}$ have the same Hilbert function and thus that $r_{{\mathbb{X}}^{\prime}}=r_{\mathbb{X}}$. Now we apply the well-known fact that $I_{\mathbb{X}}$ and $I_{{\mathbb{X}}^{\prime}}$ are generated in degrees $\leq r_{\mathbb{X}}+1$ (see for instance [21, Prop. 1.1]) to conclude that $\lambda(I_{\mathbb{X}})=I_{{\mathbb{X}}^{\prime}}$. This yields $\Lambda({\mathbb{X}})={\mathbb{X}}^{\prime}$, as was to be shown. ∎
@@ 7. The Macaulay Inverse System
In the following we assume that $J$ is homogeneous ideal in $P={\mathbb{F}}_{q}[x_{1},\dots,x_{k}]$ such that $S=P/J$ is an Artinian ring. To define the Macaulay inverse system of $S$, we first recall the divided power algebra. For further details on this ring, we refer to [24, App. A].
Let $P={\mathbb{F}}_{q}[x_{1},\dots,x_{k}]$ be equipped with the standard grading, and let
${\mathcal{D}}\;=\;\underline{\mathop{\rm Hom}\nolimits}_{{\mathbb{F}}_{q}}(P,{\mathbb{F}}_{q})\;=\;{\textstyle\bigoplus\limits_{j\geq 0}}{\mathcal{D}}_{-j}\hbox{\; with\;}{\mathcal{D}}_{-j}=\mathop{\rm Hom}\nolimits_{{\mathbb{F}}_{q}}(P_{j},{\mathbb{F}}_{q})$
be its graded dual. Since the terms in $\mathbb{T}^{n}=\{x_{1}^{\alpha_{1}}\cdots x_{k}^{\alpha_{k}}\mid\alpha_{i}\geq 0\}$ form an ${\mathbb{F}}_{q}$-basis of $P$, the dual basis, i.e., the projections $\{\pi^{[\beta]}\mid\beta=(\beta_{1},\dots,\beta_{k})\in{\mathbb{N}}^{k}\}$ form an ${\mathbb{F}}_{q}$-basis of $\mathcal{D}$}. Here we have $\pi^{[\beta]}(x^{\alpha})=\delta_{\alpha\beta}$ for $x^{\alpha}=x_{1}^{\alpha_{1}}\cdots x_{k}^{\alpha_{k}}\in\mathbb{T}^{n}$, as usual.
@@
for $\beta=(\beta_{1},\dots,\beta_{k})$ and $\gamma=(\gamma_{1},\dots,\gamma_{k})$ in ${\mathbb{N}}^{k}$. By extending this rule ${\mathbb{F}}_{q}$-linearly, we obtain an ${\mathbb{F}}_{q}$-algebra structure on $\mathcal{D}$}, and we call the resulting ring the divided power algebra in $k$ indeterminates over ${\mathbb{F}}_{q}$.
@@
${\mathbb{F}}_{q}$-linearly. This is sometimes called the action by contraction. The following facts are well-known (see [24, App. A]).
@@
Let $P={\mathbb{F}}_{q}[x_{1},\dots,x_{k}]$ and ${\mathcal{D}}=\underline{\mathop{\rm Hom}\nolimits}_{{\mathbb{F}}_{q}}(P,{\mathbb{F}}_{q})$ as above.
@@
    The multiplication defined above turns $\mathcal{D}$} into a commutative ${\mathbb{F}}_{q}$-algebra.
@@
    For every $j\geq 0$, the power products $\pi_{1}^{\beta_{1}}\cdots\pi_{k}^{\beta_{k}}$ with $\beta_{1}+\cdots+\beta_{k}=j$ form an ${\mathbb{F}}_{q}$-basis of ${\mathcal{D}}_{-j}$.
@@
    For every $j\geq 0$, the left action $\circ$ induces a non-degenerate bilinear map $\circ_{j}:\;P_{j}\times{\mathcal{D}}_{-j}\longrightarrow{\mathbb{F}}_{q}$.
@@
In general, the Macaulay inverse system of a homogeneous ideal of $P$ is not finitely generated. However, if $S=P/J$ is an Artinian ring, as we assumed above, then $J^{\perp}$ is indeed a finitely generated graded $P$-submodule of $\mathcal{D}$}. The most famous result about Macaulay inverse systems is the following theorem of F.S. Macaulay (see [35, Ch. IV]; for a proof in the modern language of algebra, see [24, Lemma 2.14]).
@@
The general linear group $G=\mathop{\rm GL}\nolimits_{k}({\mathbb{F}}_{q})$ operates on $P={\mathbb{F}}_{q}[x_{1},\dots,x_{k}]$ via homogeneous linear changes of coordinates, i.e., the map $\ast:\;G\times P\longrightarrow P$ is given by $A\ast f(x_{1},\dots,x_{n})=f(a_{11}x_{1}+\cdots+a_{1k}x_{k},\dots,a_{k1}x_{1}+\cdots+a_{kk}x_{k})$ for $A=(a_{ij})\in G$ and $f\in P$. This group action defines a dual group action on $\mathcal{D}$} in the natural way, namely as follows. Given a matrix $A\in G$, we let $\tau_{A}:\;P\longrightarrow P$ be given by $\tau_{A}(f)=A\ast f$. Then the operation of $A$ on $\mathcal{D}$} is defined by
@@
Let $J$ be a homogeneous ideal in $P={\mathbb{F}}_{q}[x_{1},\dots,x_{k}]$ such that $S=P/J$ is an Artinian Gorenstein ring. Let ${\varphi}\in{\mathcal{D}}_{-\sigma_{S}}$ be such that $J^{\perp}=\langle{\varphi}\rangle$.
@@
    Given a matrix $A\in\mathop{\rm GL}\nolimits_{k}({\mathbb{F}}_{q})$, let $J^{\prime}=A\ast J=\{A\ast f\mid f\in J\}$. Then we have $(J^{\prime})^{\perp}=\langle A\ast{\varphi}\rangle$.
@@
    Given $j\geq 1$, a matrix $A\in\mathop{\rm GL}\nolimits_{k}({\mathbb{F}}_{q})$, and an element ${\varphi}\in{\mathcal{D}}_{-j}\setminus\{0\}$, we have $\mathop{\rm Ann}\nolimits_{P}(A\ast{\varphi})=A\ast\mathop{\rm Ann}\nolimits_{P}({\varphi})$.
@@
First we show (a). Using the fact that the operation $\circ:\;P\times{\mathcal{D}}\longrightarrow{\mathcal{D}}$ is $\mathop{\rm GL}\nolimits_{k}({\mathbb{F}}_{q})$-equivariant (see [24, Prop. A.3]), we calculate
@@
Using this proposition, we can now return to our goal of reformulating the PSE problem. Recall that, for a finite set of points $\mathbb{X}$} in ${\mathbb{P}}^{k-1}$, the doubling $D_{\mathbb{X}}=R/{\mathfrak{J}}_{R/{\mathbb{F}}_{q}[\ell]}$ of $\mathbb{X}$} is an Artinian Gorenstein ${\mathbb{F}}_{q}$-algebra with socle degree $2r_{\mathbb{X}}-1$. We denote its Macaulay inverse system by ${\mathcal{I}}_{\mathbb{X}}$ and note that it is of the form ${\mathcal{I}}_{\mathbb{X}}=\langle\varPhi_{\mathbb{X}}\rangle$ with a homogeneous element $\varPhi_{\mathbb{X}}\in{\mathcal{D}}_{-2r_{\mathbb{X}}+1}={\mathbb{F}}_{q}[\pi_{1},\dots,\pi_{k}]_{-2r_{\mathbb{X}}+1}$. If we consider $\pi_{1},\dots,\pi_{k}$ as (standard graded) indeterminates, we can view $\varPhi_{\mathbb{X}}$ as a homogeneous polynomial of degree $2r_{\mathbb{X}}-1$ in ${\mathbb{F}}_{q}[\pi_{1},\dots,\pi_{k}]$. It is called the Macaulay inverse polynomial of the set of points $\mathbb{X}$}. The polynomial $\varPhi_{\mathbb{X}}$ can be found in the following way.
@@
    By [28, Cor. 1.11], the degree $2r_{\mathbb{X}}-1$ homogeneous component of the canonical ideal ${\mathfrak{J}}_{R/{\mathbb{F}}_{q}[\ell]}$ consists of all elements $\ell^{r_{\mathbb{X}}-1}\,(c_{1}f_{1}+\cdots+c_{n}f_{n})$ such that $c_{1},\dots,c_{n}\in{\mathbb{F}}_{q}$ and $c_{1}+\cdots+c_{n}=0$. In other words, we have
    $({\mathfrak{J}}_{R/{\mathbb{F}}_{q}[\ell]})_{2r_{\mathbb{X}}-1}\;=\;{\mathbb{F}}_{q}\cdot\ell^{r_{\mathbb{X}}-1}\,(f_{2}-f_{1})\oplus\cdots\oplus{\mathbb{F}}_{q}\cdot\ell^{r_{\mathbb{X}}-1}\,(f_{n}-f_{n-1})$
    Hence the residue class of $\ell^{r_{\mathbb{X}}-1}\,f_{i}$ is an ${\mathbb{F}}_{q}$-basis of $(D_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$ for every $i\in\{1,\dots,n\}$.
@@
    Next we let $\widehat{{\mathfrak{J}}}_{\mathbb{X}}$ be a preimage of ${\mathfrak{J}}_{R/{\mathbb{F}}_{q}[\ell]}$ in $P$, and let $\hat{f_{i}}\in P_{r_{\mathbb{X}}}$ be a preimage of $f_{i}$ for $i=1,\dots,n$. Thus the above implies
    $(\widehat{{\mathfrak{J}}}_{\mathbb{X}})_{2r_{\mathbb{X}}-1}\;=\;(I_{\mathbb{X}})_{2r_{\mathbb{X}}-1}\;\oplus\;{\mathbb{F}}_{q}\cdot L^{r_{\mathbb{X}}-1}\,(\hat{f}_{2}-\hat{f}_{1})\oplus\cdots\oplus{\mathbb{F}}_{q}\cdot L^{r_{\mathbb{X}}-1}\,(\hat{f}_{n}-\hat{f}_{n-1})$
    Therefore the projection $\varPhi_{\mathbb{X}}:P_{2r_{\mathbb{X}}-1}\longrightarrow{\mathbb{F}}_{q}$ to $L^{r_{\mathbb{X}}-1}\,\hat{f}_{1}$ along $(\widehat{{\mathfrak{J}}}_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$ is an ${\mathbb{F}}_{q}$-basis of $({\mathcal{I}}_{\mathbb{X}})_{-2r_{\mathbb{X}}+1}$, i.e., it is the Macaulay inverse polynomial of $\mathbb{X}$}. Notice that $\varPhi_{\mathbb{X}}$ is unique up to a scalar from ${\mathbb{F}}_{q}$ which depends on the choice of a representative in $P$ of an ${\mathbb{F}}_{q}$-basis of $(D_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$.
@@
    To get $\varPhi_{\mathbb{X}}$ as a homogeneous polynomial of degree $2r_{\mathbb{X}}-1$ in ${\mathbb{F}}_{q}[\pi_{1},\dots,\pi_{k}]$, we may proceed as follows. By multiplying the elements of a homogeneous reduced Gröbner basis of $I_{\mathbb{X}}$ with terms of the complementary degree, we get a system of generators of the ${\mathbb{F}}_{q}$-vector space $(I_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$. Then we append the elements $\{L^{r_{\mathbb{X}}-1}\,(\hat{f}_{i}-\hat{f}_{i})\mid i=2,\dots,n\}$ and interreduce the resulting list of polynomials ${\mathbb{F}}_{q}$-linearly. The result is an ${\mathbb{F}}_{q}$-basis of the 1-codimensional vector subspace $(\widehat{{\mathfrak{J}}}_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$ of $P_{2r_{\mathbb{X}}-1}$. Lastly, we use the bilinear pairing $\circ_{2r_{\mathbb{X}}-1}:\;P_{2r_{\mathbb{X}}-1}\times{\mathcal{D}}_{-2r_{\mathbb{X}}+1}\longrightarrow{\mathbb{F}}_{q}$ defined before Proposition 7.1 to compute a basis $\{\varPhi_{\mathbb{X}}\}$ of the dual vector space.
Notice that, unfortunately, the method given in part (c) of this remark has exponential complexity, since $\dim_{{\mathbb{F}}_{q}}(\widehat{{\mathfrak{J}}}_{\mathbb{X}})_{2r_{\mathbb{X}}-1}=\binom{2r_{\mathbb{X}}+k-2}{k-1}-1$. The following result improves this method substantially.
@@
First we show (a). Let $\ell$ be the residue class of $L$ in $R$. The formula for $({\mathfrak{J}}_{R/{\mathbb{F}}_{q}[\ell]})_{2r_{\mathbb{X}}-1}$ given in part (a) of the preceding remark shows that all polynomials $g\in(\widehat{{\mathfrak{J}}}_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$ satisfy $g(p_{1})+\cdots+g(p_{n})=0$. On the other hand, starting from $\mathop{\rm HF}\nolimits_{D_{\mathbb{X}}}(2r_{\mathbb{X}}-1)=1$, we see that $(\widehat{{\mathfrak{J}}}_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$ has codimension 1 in $P_{2r_{\mathbb{X}}-1}$. This implies the claim.
Now we prove (b). Let $g=\sum_{\alpha}c_{\alpha}\,x_{1}^{\alpha_{1}}\cdots x_{k}^{\alpha_{k}}$ be an element of $(\widehat{{\mathfrak{J}}}_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$, where $c_{\alpha}\in{\mathbb{F}}_{q}$ and the sum extends over all $\alpha=(\alpha_{1},\dots,\alpha_{k})\in{\mathbb{N}}^{k}$ such that $|\alpha|=2r_{\mathbb{X}}-1$. We write $\varPhi_{\mathbb{X}}=\sum_{\alpha}b_{\alpha}\,\pi_{1}^{\alpha_{1}}\cdots\pi_{k}^{\alpha_{k}}$ with $b_{\alpha}\in{\mathbb{F}}_{q}$. By construction of the Macaulay inverse polynomial, we know that $g\circ\varPhi_{\mathbb{X}}=\sum_{\alpha}c_{\alpha}b_{\alpha}=0$ for all $g\in(\widehat{{\mathfrak{J}}}_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$. The formula in part (a) yields

### v2

@@ 6. Doublings of Sets of Points
For a proof of (a), see [31, Kor. 6.13] or [10, Prop. 3.3.18]. Claim (b) follows from the Hilbert function of ${\mathfrak{J}}_{R/{\mathbb{F}}_{q}[\ell]}$ given in Section 5, and (c) follows immediately from (b). ∎
For further properties of doublings, see for instance [34, Sect. 2.5]. In the case of self-associated sets of points, we obtain the following corollary.
@@
The goal of this section is to reformulate the Point Set Equivalence Problem in several ways. A linear change of coordinates $\Lambda:\;{\mathbb{P}}^{k-1}\longrightarrow{\mathbb{P}}^{k-1}$ induces homomorphisms of several algebraic objects considered above in the following ways.
@@
$\varPhi^{\prime}(\lambda_{\omega}({\varphi}))(g)\;=\;c_{1}(\ell^{\prime})^{d}f^{\prime}_{1}+\cdots+c_{n}(\ell^{\prime})^{d}f^{\prime}_{n}\;=\;\lambda_{R}(\ell^{d}(c_{1}f_{1}+\cdots+c_{n}f_{n}))\;=\;\lambda_{R}(g)$
@@
Finally, we prove that (2) implies (1). We use $i=2$ and note that the initial degree of $\ell^{2}\,{\mathfrak{J}}_{R/{\mathbb{F}}_{q}[\ell]}$ is $r_{\mathbb{X}}+2$. Therefore the hypothesis that $\lambda_{D,2}$ is an isomorphism implies that $(\lambda_{R})_{d}:\;R_{d}\longrightarrow R^{\prime}_{d}$ is an isomorphism of ${\mathbb{F}}_{q}$-vector spaces for every $d\leq r_{\mathbb{X}}+1$. Consequently, we have $\lambda((I_{\mathbb{X}})_{d})=(I_{{\mathbb{X}}^{\prime}})_{d}$ for every $d\leq r_{\mathbb{X}}+1$. In particular, we see that $\mathbb{X}$} and $\mathbb{X}$$^{\prime}$ have the same Hilbert function and thus that $r_{{\mathbb{X}}^{\prime}}=r_{\mathbb{X}}$. Now we apply the well-known fact that $I_{\mathbb{X}}$ and $I_{{\mathbb{X}}^{\prime}}$ are generated in degrees $\leq r_{\mathbb{X}}+1$ (see for instance [26, Prop. 1.1]) to conclude that $\lambda(I_{\mathbb{X}})=I_{{\mathbb{X}}^{\prime}}$. This yields $\Lambda({\mathbb{X}})={\mathbb{X}}^{\prime}$, as was to be shown. ∎
@@ 7. The Macaulay Inverse System
The material in this section is valid over an arbitrary base field $K$. Later on, we will be mostly interested in the case of a finite field $K={\mathbb{F}}_{q}$. In the following we assume that $J$ is homogeneous ideal in $P=K[x_{1},\dots,x_{k}]$ such that $S=P/J$ is an Artinian ring. To define the Macaulay inverse system of $S$, we first recall the divided power algebra. For further details on this ring, we refer to [32, App. A].
Let $P=K[x_{1},\dots,x_{k}]$ be equipped with the standard grading, and let
${\mathcal{D}}\;=\;\underline{\mathop{\rm Hom}\nolimits}_{K}(P,K)\;=\;{\textstyle\bigoplus\limits_{j\geq 0}}{\mathcal{D}}_{-j}\hbox{\; with\;}{\mathcal{D}}_{-j}=\mathop{\rm Hom}\nolimits_{K}(P_{j},K)$
be its graded dual. Since the terms in $\mathbb{T}^{n}=\{x_{1}^{\alpha_{1}}\cdots x_{k}^{\alpha_{k}}\mid\alpha_{i}\geq 0\}$ form an $K$-basis of $P$, the dual basis, i.e., the projections $\{\pi^{[\beta]}\mid\beta=(\beta_{1},\dots,\beta_{k})\in{\mathbb{N}}^{k}\}$ form an $K$-basis of $\mathcal{D}$}. Here we have $\pi^{[\beta]}(x^{\alpha})=\delta_{\alpha\beta}$ for $x^{\alpha}=x_{1}^{\alpha_{1}}\cdots x_{k}^{\alpha_{k}}\in\mathbb{T}^{n}$, as usual.
@@
for $\beta=(\beta_{1},\dots,\beta_{k})$ and $\gamma=(\gamma_{1},\dots,\gamma_{k})$ in ${\mathbb{N}}^{k}$. By extending this rule $K$-linearly, we obtain a $K$-algebra structure on $\mathcal{D}$}, and we call the resulting ring the divided power algebra in $k$ indeterminates over $K$.
@@
$K$-linearly. This is sometimes called the action by contraction. The following facts are well-known (see [32, App. A]).
@@
Let $P=K[x_{1},\dots,x_{k}]$ and ${\mathcal{D}}=\underline{\mathop{\rm Hom}\nolimits}_{K}(P,K)$ as above.
@@
    The multiplication defined above turns $\mathcal{D}$} into a commutative $K$-algebra.
@@
    For every $j\geq 0$, the power products $\pi_{1}^{\beta_{1}}\cdots\pi_{k}^{\beta_{k}}$ with $\beta_{1}+\cdots+\beta_{k}=j$ form a $K$-basis of ${\mathcal{D}}_{-j}$.
@@
    For every $j\geq 0$, the left action $\circ$ induces a non-degenerate bilinear map $\circ_{j}:\;P_{j}\times{\mathcal{D}}_{-j}\longrightarrow K$.
@@
In general, the Macaulay inverse system of a homogeneous ideal of $P$ is not finitely generated. However, if $S=P/J$ is an Artinian ring, as we assumed above, then $J^{\perp}$ is indeed a finitely generated graded $P$-submodule of $\mathcal{D}$}. The most famous result about Macaulay inverse systems is the following theorem of F.S. Macaulay (see [43, Ch. IV]; for a proof in the modern language of algebra, see [32, Lemma 2.14]).
@@
The general linear group $G=\mathop{\rm GL}\nolimits_{k}(K)$ operates on $P=K[x_{1},\dots,x_{k}]$ via homogeneous linear changes of coordinates, i.e., the map $\ast:\;G\times P\longrightarrow P$ is given by $A\ast f(x_{1},\dots,x_{n})=f(a_{11}x_{1}+\cdots+a_{1k}x_{k},\dots,a_{k1}x_{1}+\cdots+a_{kk}x_{k})$ for $A=(a_{ij})\in G$ and $f\in P$. This group action defines a dual group action on $\mathcal{D}$} in the natural way, namely as follows. Given a matrix $A\in G$, we let $\tau_{A}:\;P\longrightarrow P$ be given by $\tau_{A}(f)=A\ast f$. Then the operation of $A$ on $\mathcal{D}$} is defined by
@@
Let $J$ be a homogeneous ideal in $P=K[x_{1},\dots,x_{k}]$ such that $S=P/J$ is an Artinian Gorenstein ring. Let ${\varphi}\in{\mathcal{D}}_{-\sigma_{S}}$ be such that $J^{\perp}=\langle{\varphi}\rangle$.
@@
    Given a matrix $A\in\mathop{\rm GL}\nolimits_{k}(K)$, let $J^{\prime}=A\ast J=\{A\ast f\mid f\in J\}$. Then we have $(J^{\prime})^{\perp}=\langle A\ast{\varphi}\rangle$.
@@
    Given $j\geq 1$, a matrix $A\in\mathop{\rm GL}\nolimits_{k}(K)$, and an element ${\varphi}\in{\mathcal{D}}_{-j}\setminus\{0\}$, we have $\mathop{\rm Ann}\nolimits_{P}(A\ast{\varphi})=A\ast\mathop{\rm Ann}\nolimits_{P}({\varphi})$.
@@
First we show (a). Using the fact that the operation $\circ:\;P\times{\mathcal{D}}\longrightarrow{\mathcal{D}}$ is $\mathop{\rm GL}\nolimits_{k}(K)$-equivariant (see [32, Prop. A.3]), we calculate
@@
Using this proposition, we can now return to our goal of reformulating the PSE problem. Recall that, for a finite set of points $\mathbb{X}$} in ${\mathbb{P}}^{k-1}$, the doubling $D_{\mathbb{X}}=R/{\mathfrak{J}}_{R/K[\ell]}$ of $\mathbb{X}$} is an Artinian Gorenstein $K$-algebra with socle degree $2r_{\mathbb{X}}-1$. We denote its Macaulay inverse system by ${\mathcal{I}}_{\mathbb{X}}$ and note that it is of the form ${\mathcal{I}}_{\mathbb{X}}=\langle\varPhi_{\mathbb{X}}\rangle$ with a homogeneous element $\varPhi_{\mathbb{X}}\in{\mathcal{D}}_{-2r_{\mathbb{X}}+1}=K[\pi_{1},\dots,\pi_{k}]_{-2r_{\mathbb{X}}+1}$. If we consider $\pi_{1},\dots,\pi_{k}$ as (standard graded) indeterminates, we can view $\varPhi_{\mathbb{X}}$ as a homogeneous polynomial of degree $2r_{\mathbb{X}}-1$ in $K[\pi_{1},\dots,\pi_{k}]$. It is called the Macaulay inverse polynomial of the set of points $\mathbb{X}$}. The polynomial $\varPhi_{\mathbb{X}}$ can be found in the following way.
@@
    By [36, Cor. 1.11], the degree $2r_{\mathbb{X}}-1$ homogeneous component of the canonical ideal ${\mathfrak{J}}_{R/K[\ell]}$ consists of all elements $\ell^{r_{\mathbb{X}}-1}\,(c_{1}f_{1}+\cdots+c_{n}f_{n})$ such that $c_{1},\dots,c_{n}\in K$ and $c_{1}+\cdots+c_{n}=0$. In other words, we have
    $({\mathfrak{J}}_{R/K[\ell]})_{2r_{\mathbb{X}}-1}\;=\;K\cdot\ell^{r_{\mathbb{X}}-1}\,(f_{2}-f_{1})\oplus\cdots\oplus K\cdot\ell^{r_{\mathbb{X}}-1}\,(f_{n}-f_{n-1})$
    Hence the residue class of $\ell^{r_{\mathbb{X}}-1}\,f_{i}$ is a $K$-basis of $(D_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$ for every $i\in\{1,\dots,n\}$.
@@
    Next we let $\widehat{{\mathfrak{J}}}_{\mathbb{X}}$ be a preimage of ${\mathfrak{J}}_{R/K[\ell]}$ in $P$, and let $\hat{f_{i}}\in P_{r_{\mathbb{X}}}$ be a preimage of $f_{i}$ for $i=1,\dots,n$. Thus the above implies
    $(\widehat{{\mathfrak{J}}}_{\mathbb{X}})_{2r_{\mathbb{X}}-1}\;=\;(I_{\mathbb{X}})_{2r_{\mathbb{X}}-1}\;\oplus\;K\cdot L^{r_{\mathbb{X}}-1}\,(\hat{f}_{2}-\hat{f}_{1})\oplus\cdots\oplus K\cdot L^{r_{\mathbb{X}}-1}\,(\hat{f}_{n}-\hat{f}_{n-1})$
    Therefore the projection $\varPhi_{\mathbb{X}}:P_{2r_{\mathbb{X}}-1}\longrightarrow K$ to $L^{r_{\mathbb{X}}-1}\,\hat{f}_{1}$ along $(\widehat{{\mathfrak{J}}}_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$ is a $K$-basis of $({\mathcal{I}}_{\mathbb{X}})_{-2r_{\mathbb{X}}+1}$, i.e., it is the Macaulay inverse polynomial of $\mathbb{X}$}. Notice that $\varPhi_{\mathbb{X}}$ is unique up to a scalar from $K$ which depends on the choice of a representative in $P$ of a $K$-basis of $(D_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$.
@@
    To get $\varPhi_{\mathbb{X}}$ as a homogeneous polynomial of degree $2r_{\mathbb{X}}-1$ in $K[\pi_{1},\dots,\pi_{k}]$, we may proceed as follows. By multiplying the elements of a homogeneous reduced Gröbner basis of $I_{\mathbb{X}}$ with terms of the complementary degree, we get a system of generators of the $K$-vector space $(I_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$. Then we append the elements $\{L^{r_{\mathbb{X}}-1}\,(\hat{f}_{i}-\hat{f}_{i})\mid i=2,\dots,n\}$ and interreduce the resulting list of polynomials $K$-linearly. The result is a $K$-basis of the 1-codimensional vector subspace $(\widehat{{\mathfrak{J}}}_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$ of $P_{2r_{\mathbb{X}}-1}$. Lastly, we use the bilinear pairing $\circ_{2r_{\mathbb{X}}-1}:\;P_{2r_{\mathbb{X}}-1}\times{\mathcal{D}}_{-2r_{\mathbb{X}}+1}\longrightarrow K$ defined before Proposition 7.1 to compute a basis $\{\varPhi_{\mathbb{X}}\}$ of the dual vector space.
Notice that, unfortunately, the method given in part (c) of this remark has exponential complexity, since $\dim_{K}(\widehat{{\mathfrak{J}}}_{\mathbb{X}})_{2r_{\mathbb{X}}-1}=\binom{2r_{\mathbb{X}}+k-2}{k-1}-1$. The following result improves this method substantially.
@@
First we show (a). Let $\ell$ be the residue class of $L$ in $R$. The formula for $({\mathfrak{J}}_{R/K[\ell]})_{2r_{\mathbb{X}}-1}$ given in part (a) of the preceding remark shows that all polynomials $g\in(\widehat{{\mathfrak{J}}}_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$ satisfy $g(p_{1})+\cdots+g(p_{n})=0$. On the other hand, starting from $\mathop{\rm HF}\nolimits_{D_{\mathbb{X}}}(2r_{\mathbb{X}}-1)=1$, we see that $(\widehat{{\mathfrak{J}}}_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$ has codimension 1 in $P_{2r_{\mathbb{X}}-1}$. This implies the claim.
Now we prove (b). Let $g=\sum_{\alpha}c_{\alpha}\,x_{1}^{\alpha_{1}}\cdots x_{k}^{\alpha_{k}}$ be an element of $(\widehat{{\mathfrak{J}}}_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$, where $c_{\alpha}\in K$ and the sum extends over all $\alpha=(\alpha_{1},\dots,\alpha_{k})\in{\mathbb{N}}^{k}$ such that $|\alpha|=2r_{\mathbb{X}}-1$. We write $\varPhi_{\mathbb{X}}=\sum_{\alpha}b_{\alpha}\,\pi_{1}^{\alpha_{1}}\cdots\pi_{k}^{\alpha_{k}}$ with $b_{\alpha}\in K$. By construction of the Macaulay inverse polynomial, we know that $g\circ\varPhi_{\mathbb{X}}=\sum_{\alpha}c_{\alpha}b_{\alpha}=0$ for all $g\in(\widehat{{\mathfrak{J}}}_{\mathbb{X}})_{2r_{\mathbb{X}}-1}$. The formula in part (a) yields